OpenStudy (anonymous):

Can anyone explains how (1/2) + ln abs (sqrt2/2) turns into (1/2)(1-ln2)?

OpenStudy (anonymous):

\[ \frac 12 +\ln\left|\frac{\sqrt{2}}{2}\right| = \frac 12 + \ln\left(\frac 24\right)^{1/2} = \frac 12 + \frac 12 \ln\left(\frac 12\right) = \frac 12 \left(1+\ln\left(2^{-1}\right)\right) \]\[ =\frac 12 \left(1-\ln\left(2\right)\right) \]

OpenStudy (anonymous):

I dont understand how the 1/2 gets in front of ln

OpenStudy (anonymous):

\[ \ln(a^b) = b\ln(a) \]

OpenStudy (anonymous):

\[ \ln(a^b) = \ln(\underbrace{a\times a\times \ldots \times a}_b) =\underbrace{ \ln(a)+\ln(a)+\ldots +\ln(a)}_b = b\ln(a) \]

OpenStudy (phi):

for some intuition, notice that ln(x^3) = ln(x * x *x) = ln(x)+ln(x)+ln(x)= 3 ln(x) so this property clearly holds for integer exponents. However, it is more general.. the property works for all exponents

OpenStudy (anonymous):

It's important to note that the reason it works for real exponents is because real exponents were initially a meaningless operation, and in defining them, they were defined such that they would obey all the properties of natural number exponents. There is no coincidence.

OpenStudy (phi):

Here is another way to see it: Let \[ x= e^a \] and \[ \ln(x) = \ln(e^a) = a \] then now raise x to a power b \[ x^b = \left(e^a\right)^b= e^{ab} \] take the ln of both sides \[ \ln\left(x^b\right) = ab \] from the second line, a = ln(x) and we get \[ \ln\left(x^b\right) = b\ \ln(x) \]